OpenStudy (anonymous):

Simplify 5√11-12√11-2√11

Vocaloid (vocaloid):

well, since all of them are sqrt(11), we can combine like terms 5-12-2=?

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

so -9√11 is the answer

Vocaloid (vocaloid):

yup!

jimthompson5910 (jim_thompson5910):

Hint When it comes to problems like this, I like to replace the square root with a variable like so \[\Large 5\sqrt{11}-12\sqrt{11}-2\sqrt{11}\] \[\Large 5{\color{red}{\sqrt{11}}}-12{\color{red}{\sqrt{11}}}-2{\color{red}{\sqrt{11}}}\] \[\Large 5{\color{red}{x}}-12{\color{red}{x}}-2{\color{red}{x}}\]

OpenStudy (anonymous):

Thank you